Find JK
JK=2x-7
KL=10
JL=3X-5
I am trying to do the two column statement and reason
I know the statement at the top is given
I think KL=JL is the statement it gives me the reason as segment addition postulate
2x-7+10=3x-5 is the statement I put subtraction postulate for reason
2x+3=3x-5 is the statement I put addition property of equality for reason
now I have subtraction propertyof equality for reason but I can't figure out the statement
8=x is the statement but I'm not sure the reason
x=8 is the statement I put division postulate for reason
please help me figure out this problem

Answer by jim_thompson5910(35256)

I am assuming that points J, K and L all lie on the same line. I'm also assuming that K is between J and L.

Number Statement Reason
1 JK+KL = JL Segment Addition Postulate
2 2x-7+10=3x-5 Substitution
3 2x+3=3x-5 Combine like terms
4 2x+3-2x=3x-5-2x Subtractive Property of Equality
5 3=x-5 Combine like terms
6 3+5=x-5+5 Additive Property of Equality
7 8=x Combine like terms
8 x=8 Symmetric Property of Equality